OpenStudy (anonymous):

What is the slope-intercept form of the equation y-7=5/2(x+4)? y-2=-(5/2) (x+2) y+7=-x+5/2 y=-(4/7)+2 y=(-(5/2)x)-3

OpenStudy (aakashsudhakar):

The first step is to note the particular forms of point-slope versus slope-intercept. POINT-SLOPE: y - y1 = m(x - x1) SLOPE-INTERCEPT: y = mx + b You have that y - 7 = (5/2)(x + 4) This way, you know that m = (5/2), x1 = -4, and y1 = 7. First, plug in the known slope (m). y = (5/2)x + b Now you need to find the value of b. The easiest way to do this is simply to plug in values for x and y and solve for b. This is very easy since we now have values for x and y: (-4, 7). 7 = (5/2)(-4) + b 7 = -10 + b b = 17 Voila, we have a value for b. Simply plug that into your slope-intercept formula, and you're done! We have now... y = (5/2)x + 17 Unfortunately, I cannot really decipher some of the answers you have given as choices, so I can't tell you with 100% accuracy which answer is the right one. I can tell you with most certainty that the slope-intercept that I've given you is probably correct.